What is the Cartesian product A × B × C, where A is the set of all airlines and B and C are both set of allcities in USA? Give an example of how this Cartesian product can be used


1
Expert's answer
2021-05-10T13:09:39-0400

The Cartesian product "A \times B \times C" consists of all the ordered triples of the form (a,b,c) where a is an airline and both b and c are cities in the United States.

One way to use the set "A \times B \times C" represents all airlines that fly from the city B to city C of the United States and "A \times B \times C" would give us information on the connectivity of the cities using the airline.

Now, we show by an example how the Cartesian product would work.

Let A = {a,b}, B = {x,y} and C = {1}.

Now the Cartesian product of A, B and C is defined as:

"A \times B \times C = {a,b} \times [{x,y} \times {1}] \\

= {a,b} \times [{x,1},{y,1}] \\

= {(a,x,1),(a,y,1),(b,x,1),(b,y,1)}"